If my point estimate is equal to 89.7 and my confidence interval is 88.7 to 90.7, then what is my margin of error?
Answer #1

Solution :

  = (Lower confidence interval + Upper confidence interval ) / 2

Sample mean = = (88.7+90.7) / 2=89.7

Margin of error = E = Upper confidence interval -   = 90.7-89.7=1

Margin of error = E =1